### v2

# Extremal contractions and non-free rational curves on Fano varieties

Abstract.
Let $X$ be a complex smooth Fano variety whose minimal anticanonical degree of non-free rational curves on $X$ is at least $\dim X-2$. We give a classification of extremal contractions of such varieties. As applications, we obtain a classification of Fano fourfolds whose pseudoindex and Picard number are greater than one and study the structure of Fano varieties with nef third exterior power of the tangent bundle.
Key words and phrases:
extremal contractions, non-free rational curves, Fano varieties
2020 Mathematics Subject Classification:
14J40, 14J45, 14E30.
The author is partially supported by JSPS KAKENHI Grant Number 21K03170.
1. Introduction
A complex smooth projective variety $X$ is called Fano if its anticanonical divisor $-K_{X}$ is ample. In both classical and modern perspectives in projective geometry and the Minimal Model Program, Fano varieties carry substantial significance. It is known that there are finitely many families of smooth Fano varieties for each dimension [39]; in particular, Fano threefolds were classified by Fano, Iskovskih, Shokurov, Fujita, Mori, Mukai and Takeuchi (see [56] and references therein). The classification becomes more challenging in dimensions four and higher; Fano fourfolds have been investigated by many authors (see some examples [70, 69, 54, 41, 63, 64, 14, 13, 25, 10, 15, 16, 11, 30]), but they are still far from a complete classification.
For a Fano $n$-fold $X$ with $n\geq 3$, we define the Fano index $i_{X}$ and the pseudoindex $\iota_{X}$ of $X$ by
$i_{X}:=\max\left\{m\in{\mathbb{Z}}\,\,\middle|\,\,-K_{X}=mL~{}\mbox{with}~{}L\in{\rm Pic}(X)\right\}$
and
$\iota_{X}:=\min\{(-K_{X})\cdot C\mid C\subset X~{}\mbox{is a rational curve}\}$
respectively. Remark that $i_{X}$ divides $\iota_{X}$. By Kobayashi-Ochiai Theorem [37], $i_{X}\leq n+1$; moreover if $i_{X}=n+1$, then $X$ is isomorphic to a projective space ${\mathbb{P}}^{n}$, and if $i_{X}=n$, then $X$ is isomorphic to a smooth quadric $Q^{n}$. If $i_{X}=n-1$, then $X$ is called a del Pezzo variety, which is completely classified by Fujita [26, 27]; if $i_{X}=n-2$, then $X$ is called a Mukai variety, which is completely classified by Mukai [49]. On the other hand, as a generalization of Kobayashi-Ochiai Theorem [37], ${\iota}_{X}\leq\dim X+1$; moreover if ${\iota}_{X}=n+1$, then $X$ is ${\mathbb{P}}^{n}$ [18], and if ${\iota}_{X}=n$, then $X$ is $Q^{n}$ [47, 21]. Unlike the case of Fano indices, a classification of Fano varieties with pseudoindices $n-1$ and $n-2$ is widely open.
The main object in this paper is a Fano varieties with Picard number greater than one and these extremal contractions. Extremal contractions play a crucial role in understanding the geometry of higher-dimensional Fano varieties. To explain our results, we introduce a new invariant of Fano varieties as follows. Let $X$ be a smooth Fano variety. A non-trivial morphism $f:{\mathbb{P}}^{1}\to X$ is called a free rational curve if $f^{\ast}T_{X}$ is nef, where $T_{X}$ is the tangent bundle of $X$. It is called a non-free rational curve otherwise. We denote by ${\rm Hom}^{\rm nf}({\mathbb{P}}^{1},X)$ the closed subscheme of the Hom scheme ${\rm Hom}({\mathbb{P}}^{1},X)$ parametrizing non-free rational curves.
Definition 1.1.
For a smooth Fano variety $X$,
$\tau_{X}:=\begin{cases}\min\bigl{\{}\mathop{\rm deg}\nolimits f^{\ast}(-K_{X})\mid[f]\in{\rm Hom}^{\rm nf}({\mathbb{P}}^{1},X)\bigr{\}}&\text{if ${\rm Hom}^{\rm nf}({\mathbb{P}}^{1},X)\neq\emptyset$,}\\ +\infty&\text{otherwise.}\end{cases}$
Our main results are the following:
Theorem 1.2.
Let $X$ be a smooth Fano variety of dimension $n\geq 4$ and $\tau_{X}\geq n-2$. Let $\varphi:X\to Y$ be a contraction of an extremal ray. If $\varphi$ is of fiber type, then $\varphi$ is one of the following:
  1. (i)
    a smooth morphism;
  2. (ii)
    a ${\mathbb{P}}^{n-3}$-fibration which is not equidimensional;
  3. (iii)
    an equidimensional $Q^{n-2}$-fibration.
Moreover any fiber of $\varphi$ has codimension at least two if $X$ is not a product of ${\mathbb{P}}^{1}$ and a variety.
Following the idea of the proof of [73], we also prove the following:
Theorem 1.3.
Let $X$ be a smooth Fano variety of dimension $n\geq 4$. Assume that $\tau_{X}\geq n-2$. If $X$ admits a birational contraction of an extremal ray, then $X$ is isomorphic to one of the following:
  1. (i)
    ${\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{n-1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{n-1}}(-1))$;
  2. (ii)
    ${\mathbb{P}}({\mathcal{O}}_{Q^{n-1}}\oplus{\mathcal{O}}_{Q^{n-1}}(-1))$;
  3. (iii)
    ${\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{n-1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{n-1}}(-2))$;
  4. (iv)
    the blow up of ${\mathbb{P}}^{n}$ along a line;
  5. (v)
    the blow up of $Q^{4}$ along a line;
  6. (vi)
    the blow up of $Q^{4}$ along a conic which is not on a plane contained in $Q^{4}$;
  7. (vii)
    ${\mathbb{P}}^{1}\times{\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{n-2}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{n-2}}(-1))$.
For a smooth Fano fourfold $X$, $\tau_{X}\geq n-2$ if and only if $\iota_{X}\geq 2$ (see Lemma 5.1). As an application of Theorem 1.2 and 1.3, we obtain a complete classification of Fano fourfolds whose pseudoindex and Picard number are greater than one:
Theorem 1.4.
Let $X$ be a smooth Fano fourfold. If $\iota_{X}\geq 2$ and $\rho_{X}\geq 2$, then one of the following holds.
  1. (i)
    $X$ is isomorphic to ${\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{3}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{3}}(-1))$.
  2. (ii)
    $X$ is isomorphic to the blow up of ${\mathbb{P}}^{4}$ along a line.
  3. (iii)
    $X$ is isomorphic to ${\mathbb{P}}^{1}\times{\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{2}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{2}}(-1))$.
  4. (iv)
    $X$ is isomorphic to a divisor on ${\mathbb{P}}^{2}\times{\mathbb{P}}^{3}$ of bidegree $(1,1)$.
  5. (v)
    $X$ is isomorphic to ${\mathbb{P}}^{2}\times{\mathbb{P}}^{2}$.
  6. (vi)
    $X$ is a Fano fourfold of $i_{X}=2$.
Remark that Fano fourfolds of index two and Picard number greater than one are completely classified in [69].
As another application of Theorem 1.2 and 1.3, we generalize all main results for varieties with nef $\wedge^{2}T_{X}$ in [67] to the case where $\wedge^{3}T_{X}$ is nef. More precisely, for a smooth Fano variety of dimension $n\geq 4$, the condition $\tau_{X}\geq n-2$ holds if $\wedge^{r}T_{X}$ is nef for $1\leq r\leq 3$ (see Lemma 6.2). Thus Theorem 1.2 and 1.3 can be regarded as a generalization of [22, Theorem 5.2], [58, Theorem 4.4] and [67, Theorem 1.7] for Fano varieties. See for details in Section 6.
2. Preliminaries
Notation and Conventions
We will make use of the standard notation as in [29], [38], [40], [43] and [44]. Along this paper, we work over the complex number field.
  • For a projective variety $X$, we define $N_{1}(X)$ (resp. $N^{1}(X)$) as the vector space of one-cycles (resp. one-cocycle) on $X$ with real coefficients, considering numerical equivalence. The dimension of $N_{1}(X)$ is referred to as the Picard number of $X$ denoted by $\rho_{X}$.
  • For a projective variety $X$, ${\rm Chow}(X)$ stands for the Chow scheme parametrizing cycles on $X$ (see [38, Section I.3] for details).
  • By a curve, we mean a projective variety of dimension one.
  • An $F$-bundle over a smooth projective variety $X$ is a smooth morphism $f:Y\to X$ between smooth projective varieties, with each fiber isomorphic to a variety $F$. We denote by $T_{f}$ the relative tangent bundle of $f:Y\to X$.
  • An $F$-fibration over a projective variety $X$ is a surjective morphism $f:Y\to X$ between projective varieties, with general fiber isomorphic to a variety $F$.
  • Given a vector bundle $E$ on a projective variety $X$, we use ${\mathbb{P}}(E)$ to represent the projectivization ${\rm Proj}({\rm Sym},E)$, and ${\mathcal{O}}_{{\mathbb{P}}(E)}(1)$ to denote the tautological line bundle on ${\mathbb{P}}(E)$.
  • ${\mathbb{P}}^{n}$ denotes projective $n$-space, while $Q^{n}$ represents a smooth quadric hypersurface in ${\mathbb{P}}^{n+1}$.
2.1. Extremal contractions
Let $X$ be a smooth projective variety, and let $\varphi:X\to Y$ be a contraction of a $K_{X}$-negative extremal ray $R$, referred to as an elementary contraction. Every general fiber of $\varphi$ is a smooth Fano variety. The length of the ray $R$ is defined as
[…]
A rational curve $C$ on $X$ is termed a minimal extremal rational curve of $R$ if $[C]\in R$ and $-K_{X}\cdot C=\ell(R)$. The structure of a contraction of a $K_{X}$-negative extremal ray $R$ is constrained by the following inequality, which relates the dimensions of the exceptional divisor and the fiber, as well as the length of $R$:
Proposition 2.1 (Ionescu-Wiśniewski inequality [33, Theorem 0.4], [71, Theorem 1.1] ).
Let $X$ be a smooth projective variety, and let $\varphi:X\to Y$ be a contraction of a $K_{X}$-negative extremal ray $R$, with $E$ representing its exceptional locus. Additionally, consider $F$ as an irreducible component of a non-trivial fiber of $\varphi$. Then
$\dim E+\dim F\geq\dim X+\ell(R)-1.$
2.2. Deformation theory of rational curves
To explore the structure of a uniruled variety, it is customary to investigate families of rational curves. Let us recall the space of rational curves $\mathop{\rm RatCurves}\nolimits^{n}(X)$, constructed as the normalization of the subscheme of the Chow scheme $\mathop{\rm Chow}\nolimits(X)$ that parametrizes rational curves on $X$. An irreducible component of $\mathop{\rm RatCurves}\nolimits^{n}(X)$ is referred to as a family of rational curves ${\mathcal{M}}$ on $X$. This family ${\mathcal{M}}$ is equipped with a ${\mathbb{P}}^{1}$-bundle $p:{\mathcal{U}}\to{\mathcal{M}}$ and an evaluation morphism $q:{\mathcal{U}}\to X$, which is described in detail in [38, Section II.2]. A rational curve parametrized by ${\mathcal{M}}$ is called an ${\mathcal{M}}$-curve. Since any ${\mathcal{M}}$-curve is numerically equivalent to each other, the family ${\mathcal{M}}$ determines a numerical class, we denote it by $[{\mathcal{M}}]\in N_{1}(X)$. The anticanonical degree of the family ${\mathcal{M}}$ refers to the intersection number denoted as $\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}$, which is calculated as $-K_{X}\cdot C$, where $[C]$ represents any curve belonging to the family ${\mathcal{M}}$. A half line generated by the class $[M]\in N_{1}(X)$ is denoted by ${\mathbb{R}}_{\geq 0}[{\mathcal{M}}]$. The union of all ${\mathcal{M}}$-curves is denoted by ${\rm Locus}({\mathcal{M}})$. For a point $x\in X$, the normalization of $p(q^{-1}(x))$ is denoted by ${\mathcal{M}}_{x}$, and we use ${\rm Locus}({\mathcal{M}}_{x})$ to refer to the union of all ${\mathcal{M}}_{x}$-curves.
Definition 2.2.
Under the above notation:
  1. (i)
    ${\mathcal{M}}$ is a dominating family (resp. covering family) if the evaluation morphism $q:{\mathcal{U}}\to X$ is dominant (resp. surjective).
  2. (ii)
    ${\mathcal{M}}$ is locally unsplit if, for a general point $x\in{\rm Locus}({\mathcal{M}})$, ${\mathcal{M}}_{x}$ is proper.
  3. (iii)
    ${\mathcal{M}}$ is unsplit if ${\mathcal{M}}$ is proper.
For instance, a family of rational curves ${\mathcal{M}}$ is locally unsplit if it is a dominating family with minimal degree concerning some ample line bundle on $X$. A family of rational curves ${\mathcal{M}}$ is a dominating family if and only if there exists a free ${\mathcal{M}}$-curve (see, for instance, [38, IV Theorem 1.9]). By definition, an unsplit family is locally unsplit, and an unsplit dominating family is a covering family. Mori’s bend and break [48, Theorem 4] tells us that a locally unsplit family of rational curves ${\mathcal{M}}$ satisfies $\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}\leq\dim X+1$. We often use the following results:
Lemma 2.3 ([68, Lemma 2.7]).
Let $X$ be a smooth projective variety and ${\mathcal{M}}\subset\mathop{\rm RatCurves}\nolimits^{n}(X)$ a family of rational curves. If ${\mathcal{M}}$ is not proper, then there exists a rational $1$-cycle $Z=\sum_{i=1}^{s}a_{i}Z_{i}$ satisfying the following conditions:
  1. (i)
    $Z$ is algebraically equivalent to ${\mathcal{M}}$-curves, where each $a_{i}$ is a positive integer and each $Z_{i}$ is a rational curve;
  2. (ii)
    $\sum_{i=1}^{s}a_{i}\geq 2$.
We term this rational $1$-cycle $Z=\sum_{i=1}^{s}a_{i}Z_{i}$ as a degeneration of ${\mathcal{M}}$-curves.
Proposition 2.4 ([38, IV Corollary 2.6]).
Let $X$ be a smooth projective variety and ${\mathcal{M}}$ a locally unsplit family of rational curves on $X$. For a general point $x\in{\rm Locus}({\mathcal{M}})$,
$\dim{\rm Locus}({\mathcal{M}}_{x})\geq\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}+\mathop{\rm codim}\nolimits_{X}{\rm Locus}({\mathcal{M}})-1.$
Moreover if ${\mathcal{M}}$ is unsplit, the above inequality holds for any point $x\in{\rm Locus}({\mathcal{M}})$.
For a smooth projective variety $X$, let ${\mathcal{M}}$ be an unsplit covering family of rational curves on $X$. Two points $x_{1},x_{2}\in X$ are ${\mathcal{M}}$-equivalent if they can be connected by a connected chain of ${\mathcal{M}}$-curves. By [8, 39] (see also [20, Chapter 5] and [38, IV Theorem 4.16]), there exists a proper surjective morphism $\varphi:X^{0}\to Y$ from a nonempty open subset $X^{0}\subset X$ onto a normal variety $Y$ whose fibers are ${\mathcal{M}}$-equivalent classes. We call $\varphi:X^{0}\to Y$ the rationally chain connected fibration with respect to ${\mathcal{M}}$ (RCC-fibration with respect to ${\mathcal{M}}$ for short). When $X^{0}$ coincides with the whole space $X$, then we say that $\varphi:X=X^{0}\to Y$ is the geometric quotient for ${\mathcal{M}}$.
Theorem 2.5 (cf. [22, Theorem 5.2], [58, Theorem 4.4], [35, Theorem 2.2, 2.3]).
Let $X$ be a smooth projective variety and ${\mathcal{M}}$ its unsplit covering family of rational curves. If any ${\mathcal{M}}$-curve is free, then ${\mathbb{R}}_{\geq 0}[{\mathcal{M}}]$ is an extremal ray of $\overline{{\mathop{\rm NE}\nolimits}}(X)$ and the associated contraction is a smooth geometric quotient $\varphi:X\to Y$ for ${\mathcal{M}}$.
Proof.
This follows from the same proof as in [36, Proposition 4.4].
Theorem 2.6.
Let $X$ be a smooth projective variety and ${\mathcal{M}}$ its unsplit covering family of rational curves. If $\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}\geq\dim X-2$, then ${\mathbb{R}}_{\geq 0}[{\mathcal{M}}]$ is an extremal ray of $\overline{{\mathop{\rm NE}\nolimits}}(X)$ and the associated contraction is a geometric quotient $\varphi:X\to Y$ for ${\mathcal{M}}$.
Proof.
This follows from Proposition 2.4 and [7, Theorem 2].
Remark 2.7.
Instead of the family of rational curves, we often deal with the family of $1$-cycles, like [7]. Theorem 2.6 also holds for an irreducible closed subset of ${\rm Chow}(X)$ which parametrizes $1$-cycles. See [7] for details.
3. Contractions of fiber type
Proposition 3.1.
Let $X$ be a smooth Fano variety of dimension $n$ and $\iota_{X}\geq 2$. Assume that there exists a contraction $\varphi:X\to{\mathbb{P}}^{1}$. Then $X$ is a product of ${\mathbb{P}}^{1}$ and a variety $Z$.
Proof.
By [28], there exists a minimal section ~$\tilde{\ell}\subset X$ of $\varphi$, that is, ~$\tilde{\ell}\subset X$ is a section of $\varphi$ whose anticanonical degree is minimal among such sections. Let us take a family of rational curves ${\mathcal{M}}\subset\mathop{\rm RatCurves}\nolimits^{n}{(X)}$ containing $[\tilde{\ell}]$. Then it follows from Lemma 2.3 that ${\mathcal{M}}$ is unsplit. We first claim that $\dim{\rm Locus}({\mathcal{M}}_{x})=1$ for any $x\in{\rm Locus}({\mathcal{M}})$. To prove this, assume the contrary, that is, $\dim{\rm Locus}({\mathcal{M}}_{x_{0}})\geq 2$ for some $x_{0}\in{\rm Locus}({\mathcal{M}})$. Then we may find a point $o\in{\mathbb{P}}^{1}$ such that $\varphi^{-1}(o)\cap{\rm Locus}({\mathcal{M}}_{x_{0}})\neq\emptyset$. Then we have
$\dim\left(\varphi^{-1}(o)\cap{\rm Locus}({\mathcal{M}}_{x_{0}})\right)\geq\dim\varphi^{-1}(o)+\dim{\rm Locus}({\mathcal{M}}_{x_{0}})-n\geq 1.$
Thus there exits a curve $C\subset\varphi^{-1}(o)\cap{\rm Locus}({\mathcal{M}}_{x_{0}})$. Applying [38, II, Corollary 4.21], we obtain $[C]\in{\mathbb{R}}_{\geq 0}[{\mathcal{M}}]$. This implies that any ${\mathcal{M}}$-curve is contracted by $\varphi$; this is a contradiction.
Let $x$ be any point on ${\rm Locus}({\mathcal{M}})$. By Proposition 2.4, we have
$\dim{\rm Locus}({\mathcal{M}}_{x})\geq\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}+\mathop{\rm codim}\nolimits_{X}{\rm Locus}({\mathcal{M}})-1\geq 1+\mathop{\rm codim}\nolimits_{X}{\rm Locus}({\mathcal{M}})..$
Thus we see that ${\mathcal{M}}$ is an unsplit covering family. Denote by $\psi:X\dashrightarrow Z$ the RCC-fibration with respect to ${\mathcal{M}}$. By the same argument as in the above claim that $\dim{\rm Locus}({\mathcal{M}}_{x})=1$ for any $x\in{\rm Locus}({\mathcal{M}})$, we may show that $\dim{\rm ChLocus}_{m}({\mathcal{M}})_{x}=1$ for any $x\in X$. Thus [7, Proposition 1] tells us that $\psi$ is a morphism. This yields that $\psi$ is the contraction associated to the ray ${\mathbb{R}}_{\geq 0}[{\mathcal{M}}]$. Since $\dim{\rm ChLocus}_{m}({\mathcal{M}})_{x}=1$ for any $x\in X$, we see that $\psi$ is a conic bundle; moreover any fiber of $\psi$ is numerically equivalent and any ${\mathcal{M}}$-curve is a section of $\varphi$. This concludes that $\psi$ is a ${\mathbb{P}}^{1}$-bundle. By construction, we see that $\varphi\times\psi:X\to{\mathbb{P}}^{1}\times Z$ is bijective; thus the Zariski main theorem concludes that $X\cong{\mathbb{P}}^{1}\times Z$.
Proof of Theorem 1.2.
Let $\varphi:X\to Y$ be a contraction of an extremal ray $R$ as in Theorem 1.2. If $\dim Y=1$, then it follows from Proposition 3.1 that $\varphi$ is a smooth morphism; since $Y$ is isomorphic to ${\mathbb{P}}^{1}$, Proposition 3.1 implies that $X$ is isomorphic to a product of ${\mathbb{P}}^{1}$ and a variety. Thus we assume that $\dim Y\geq 2$. General fibers of $\varphi$ are smooth Fano varieties; so there exists a dominating family ${\mathcal{M}}\subset\mathop{\rm RatCurves}\nolimits^{n}(X)$ such that any ${\mathcal{M}}$-curve is contracted by $\varphi$; replacing if necessary, we may assume that the anticanonical degree of ${\mathcal{M}}$ is minimal among such families; then ${\mathcal{M}}$ is locally unsplit. By Mori’s bend and break, we have $\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}\leq n+1$. If $\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}\geq n$, then Proposition 2.4 yields that $\dim{\rm Locus}({\mathcal{M}}_{x})\geq n-1$ for a general point $x\in X$. This contradicts to our assumption that $\dim Y\geq 2$. Thus we obtain the following inequality:
$n-1\geq\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}\geq\ell(R).$
Then we see that ${\mathcal{M}}$ is unsplit; in fact, if it is not, by Lemma 2.3 there exists a degeneration of ${\mathcal{M}}$-curves $Z=\sum_{i=1}^{s}a_{i}Z_{i}$ such that $\sum_{i=1}^{s}a_{i}>1$. By the minimality of $\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}$, each $Z_{i}$ is not free; thus by our assumption that $\tau_{X}\geq n-2$, we have
$n-1\geq\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}=-K_{X}\cdot Z\geq 2(n-2).$
This is a contradiction.
Assume that any minimal extremal rational curve of the ray $R$ is free. Then ${\mathcal{M}}$ consists of these minimal extremal rational curves and all ${\mathcal{M}}$-curves are free; then Theorem 2.5 implies that $\varphi:X\to Y$ is a smooth morphism. So assume that there exists a minimal extremal rational curve of the ray $R$ which is not free. Then by our assumption we have $\ell(R)\geq n-2$ and we obtain the inequality
$n-1\geq\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}\geq\ell(R)\geq n-2.$
For any $x\in X$, again by Proposition 2.4, we have
$\dim{\rm Locus}({\mathcal{M}}_{x})\geq\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}-1\geq n-3.$
By Theorem 2.6, ${\mathbb{R}}_{\geq 0}[{\mathcal{M}}]$ is an extremal ray and the associated contraction is a geometric quotient $\varphi:X\to Y$ for ${\mathcal{M}}$. Let us denote by $B$ the jumping locus for fiber-dimension of $\varphi$:
$B:=\{x\in X\mid\dim\varphi^{-1}(\varphi(x))>\dim X-\dim Y\}.$
By [7, Proposition 1], we obtain $\dim B\leq n-2$. Then we claim the following:
Claim 3.2.
$\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}=\ell(R)$.
Proof of the Claim 3.2.
Assume that $\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}>\ell(R)$. Then $\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}=n-1$ and $\ell(R)=n-2$. Moreover it turns out that, for any point $x\in X$,
$n-2\geq\dim X-\dim Y\geq\dim{\rm Locus}({\mathcal{M}}_{x})\geq\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}-1=n-2.$
This yields $(\dim Y,\dim{\rm Locus}({\mathcal{M}}_{x}))=(2,n-2)$. Since $\dim{\rm Locus}({\mathcal{M}}_{x})=n-2=\dim X-\dim Y$, [18] implies that general fibers of $\varphi$ are isomorphic to ${\mathbb{P}}^{n-2}$. Moreover $\varphi$ is equidimensional, because $\dim B\leq n-2$. Since $\varphi$ is an equidimensional morphism between smooth varieties, it is flat. Following the idea of the proof of [31, Proposition 3.1], we shall prove that $\varphi$ is a projective bundle; then we get a contradiction to our assumption $\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}>\ell(R)$.
To prove $\varphi$ is a projective bundle, assume the contrary. Then $\varphi$ is not smooth. We denote by $S_{\varphi}$ the locus of points of $Y$ over which $\varphi$ is not smooth and by $R_{\varphi}$ the locus of points of $Y$ over which the fibers of $\varphi$ are reducible. According to the rigidity of the projective space [57, Main Theorem] (see also [32, Theorem 1’]), we remark that $\varphi|_{X\setminus\varphi^{-1}(S_{\varphi})}:X\setminus\varphi^{-1}(S_{\varphi})\to Y\setminus S_{\varphi}$ is a ${\mathbb{P}}^{n-2}$-bundle. It follows from [5, Theorem 4] that $S_{\varphi}$ is of pure codimension one and $S_{\varphi}=\overline{R_{\varphi}}$. Since $\dim Y=2$, the locus $S_{\varphi}$ is a union of curves.
Let $C_{1}$ be a minimal extremal rational curve of the ray $R$. Taking a family of rational curves ${\mathcal{M}}^{\prime}\subset\mathop{\rm RatCurves}\nolimits^{n}(X)$ containing $[C_{1}]$, the minimality of the anticanonical degree $-K_{X}\cdot C_{1}$ and Lemma 2.3 imply that ${\mathcal{M}}^{\prime}$ is unsplit. By Proposition 2.4, for any $x\in{\rm Locus}({\mathcal{M}}^{\prime})$, we obtain the inequality
(1) $1\displaystyle 2n-2=\dim X+\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}^{\prime}\leq\dim{\rm Locus}({\mathcal{M}}^{\prime})+\dim{\rm Locus}({\mathcal{M}}_{x}^{\prime})+$ 1 .
Since any ${\mathcal{M}}^{\prime}$-curve is not free, ${\rm Locus}({\mathcal{M}}^{\prime})$ does not coincide with the whole space $X$. Moreover, for any $x\in{\rm Locus}({\mathcal{M}}^{\prime})$, ${\rm Locus}({\mathcal{M}}_{x}^{\prime})$ is contained in the fiber $\varphi^{-1}(\varphi(x))$. Hence the right hand side of the inequality $\rm(1)$ is at most
$(n-1)+(n-2)+1=2n-2.$
As a consequence, we see that
$\dim{\rm Locus}({\mathcal{M}}^{\prime})=n-1\,\,\,\,\,\mbox{and}\,\,\,\,\,\dim{\rm Locus}({\mathcal{M}}_{x}^{\prime})=n-2.$
Since ${\mathcal{M}}^{\prime}$ is unsplit, Mori’s bend and break tells us that the evaluation map from the universal family of ${\mathcal{M}}_{x}^{\prime}$ onto ${\rm Locus}({\mathcal{M}}_{x}^{\prime})$ is generically finite; this yields that $\dim{\mathcal{M}}_{x}^{\prime}=n-3$ for any $x\in{\rm Locus}({\mathcal{M}}^{\prime})$. Consider the universal family $p:{\mathcal{U}}^{\prime}\to{\mathcal{M}}^{\prime}$ associated to ${\mathcal{M}}^{\prime}$ and its evaluation morphism $q:{\mathcal{U}}^{\prime}\to{\rm Locus}({\mathcal{M}}^{\prime})$; by dimension count, we obtain
$\dim{\mathcal{M}}^{\prime}=\dim{\mathcal{U}}^{\prime}-1=\dim{\mathcal{M}}_{x}^{\prime}+\dim{\rm Locus}({\mathcal{M}}^{\prime})-1=(n-3)+(n-1)-1=2n-5.$
We see that ${\rm Locus}({\mathcal{M}}^{\prime})$ is contained in $\varphi^{-1}(S_{\varphi})$; since we have
$\dim{\rm Locus}({\mathcal{M}}^{\prime})=\dim\varphi^{-1}(S_{\varphi})=n-1,$
${\rm Locus}({\mathcal{M}}^{\prime})$ is an irreducible component of $\varphi^{-1}(S_{\varphi})$. Moreover $\varphi({\rm Locus}({\mathcal{M}}^{\prime}))$ coincides with an irreducible component of $S_{\varphi}=\overline{R_{\varphi}}$; thus we may find a point $x_{0}\in{\rm Locus}({\mathcal{M}}^{\prime})$ such that $\varphi(x_{0})$ is in $R_{\varphi}$. Set $F_{0}:=\varphi^{-1}(\varphi(x_{0}))$. Denoting by $D_{1},\ldots,D_{m}~{}(m>1)$ irreducible components of the fiber $(F_{0})_{\rm red}$ with reduced structure, there exist positive integers $a_{i}$’s such that the cycle $[F_{0}]$ is algebraically equivalent to $\sum_{i=1}^{m}a_{i}[D_{i}]$ with $\sum_{i=1}^{m}a_{i}>1$. Set $D_{1}^{o}:=\left(D_{1}\setminus\bigcup_{i\neq 1}D_{i}\right)$, which is a nonempty open subset of $D_{1}$. Without loss of generality, we may assume that $D_{1}\subset{\rm Locus}({\mathcal{M}}^{\prime})$. For any point $x\in D_{1}^{o}$, any ${\mathcal{M}}_{x}^{\prime}$-curve is contained in $F_{0}$ and passes through the point $x\in X$; this implies that ${\rm Locus}({\mathcal{M}}_{x}^{\prime})$ is contained in $D_{1}$; since ${\rm Locus}({\mathcal{M}}_{x}^{\prime})$ and $D_{1}$ have the same dimension and $D_{1}$ is irreducible, we see that ${\rm Locus}({\mathcal{M}}_{x}^{\prime})=D_{1}$. By a dimension count, we obtain
$\dim p(q^{-1}(D_{1}^{o}))=\dim q^{-1}(D_{1}^{o})-1=\dim{\mathcal{M}}_{x}^{\prime}+\dim D_{1}^{o}-1=2n-6.$
This implies that there exists a family of rational curves ${\mathcal{N}}\subset\mathop{\rm RatCurves}\nolimits^{n}(D_{1})$ such that $\dim{\mathcal{N}}$ is at least $62n-$ 6 and $\mathop{\rm deg}\nolimits_{(-K_{X}|_{D_{1}})}{\mathcal{N}}=\ell(R)$. Denoting by $\nu:\tilde{D_{1}}\to D_{1}$ the normalization of $D_{1}$, our family ${\mathcal{N}}$ lifts to $\tilde{D_{1}}$. We denote by $\tilde{{\mathcal{N}}}$ the lifting of ${\mathcal{N}}$ to $\tilde{D_{1}}$. Then by applying [31, Theorem 2.1], we see that $D_{1}$ is isomorphic to ${\mathbb{P}}^{n-2}$ and $\tilde{{\mathcal{N}}}$ is the family of lines on ${\mathbb{P}}^{n-2}$.
Let $F$ be a general fiber of $\varphi$. By [38, I, Proposition 3.12], we obtain
$(n-1)^{n-2}$ $=$ $(-K_{F})^{n-2}=[F]\cdot(-K_{X})^{n-2}=[F_{0}]\cdot(-K_{X})^{n-2}$
$=$ $\sum_{i=1}^{m}a_{i}\left([D_{i}]\cdot(-K_{X})^{n-2}\right)>[D_{1}]\cdot(-K_{X})^{n-2}=\left(\nu^{\ast}(-K_{X})\right)^{n-2}.$
For a positive integer $b$ such that $\nu^{\ast}(-K_{X})={\mathcal{O}}_{{\mathbb{P}}^{n-2}}(b)$, the above inequality tells us that $b^{n-2}<(n-1)^{n-2}$; thus we see that $b<n-1$. On the other hand, for a line $f\subset{\mathbb{P}}^{n-2}$, we have
$n-2=\ell(R)\leq(-K_{X})\cdot\nu_{\ast}(f)=\nu^{\ast}(-K_{X})\cdot f=b<n-1.$
Hence we obtain $b=n-2$.
Recall that ${\mathcal{M}}$ is an unsplit covering family; for any point $x\in D_{1}^{o}$ there exists an ${\mathcal{M}}_{x}$-curve $C$; we see that $C$ is contained in $D_{1}$; it lifts to a curve $\tilde{C}$ in $\tilde{D_{1}}\cong{\mathbb{P}}^{n-2}$. Then we obtain the following:
$(n-2){\mathcal{O}}_{{\mathbb{P}}^{n-2}}(1)\cdot\tilde{C}=\nu^{\ast}(-K_{X})\cdot\tilde{C}=(-K_{X})\cdot{C}=n-1.$
This is a contradiction, because ${\mathcal{O}}_{{\mathbb{P}}^{n-2}}(1)\cdot\tilde{C}$ is an integer and $n\geq 4$.
Let $F$ be a smooth general fiber of $\varphi$. As we proved, we have the following inequalities:
$n-2\geq\dim F\geq n-3,$
$\quad n-1\geq\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}=\ell(R)\geq n-2.$
By [18] and [21] (see also [47]), $F$ is isomorphic to ${\mathbb{P}}^{n-2},Q^{n-2}$ or ${\mathbb{P}}^{n-3}$. Assume that $F$ is isomorphic to ${\mathbb{P}}^{n-2}$ or $Q^{n-2}$; then $\varphi$ is equidimensional, because $\dim B\leq n-2$. Thus if $F$ is isomorphic to ${\mathbb{P}}^{n-2}$, [31] yields that $\varphi$ is a ${\mathbb{P}}^{n-2}$-bundle; this contradicts to our assumption that $\varphi$ is not smooth. As a consequence, we see that $\varphi$ is an equidimensional $Q^{n-2}$-fibration. Finally, we assume that $F$ is isomorphic to ${\mathbb{P}}^{n-3}$. If $\varphi$ is equidimensional, then again by [31] $\varphi$ should be a ${\mathbb{P}}^{n-3}$-bundle; it is a contradiction because $\varphi$ is not smooth. As a consequence, $\varphi$ is a ${\mathbb{P}}^{n-3}$-fibration which is not equidimensional.
4. Contractions of birational type
By the same argument as in [73] and [67, Theorem 1.7], we obtain the following:
Theorem 4.1 ([73], [67, Theorem 1.7]).
Let $X$ be a smooth Fano variety of dimension $n=\dim X\geq 4$ and $\tau_{X}\geq n-1$. If $X$ admits a birational elementary contraction, then $X$ is isomorphic to ${\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{n-1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{n-1}}(-1))$ and $\tau_{X}=n-1$.
In this section, we prove Theorem 1.3. It was proved by Yasutake in [73] when $\wedge^{3}T_{X}$ is nef; however it seems that his proof contains a nontrivial gap (see Remark 4.14); thus, following the idea of the proof of [73], we fill the gap. Before giving the proof of Theorem 1.3, we confirm that all varieties appearing in the theorem satisfy the condition that $\tau_{X}=n-2$. To prove this, it is enough to show that $\wedge^{3}T_{X}$ is nef. In fact, combining this with Theorem 4.1 and Lemma 6.2, we see that $\tau_{X}=n-2$.
Example 4.2.
Let $X={\mathbb{P}}({\mathcal{O}}_{Q^{n-1}}\oplus{\mathcal{O}}_{Q^{n-1}}(1))\cong{\mathbb{P}}({\mathcal{O}}_{Q^{n-1}}\oplus{\mathcal{O}}_{Q^{n-1}}(-1))$ for $n\geq 4$. We denote by ${\mathcal{O}}_{X}(1)$ the tautological line bundle of $X={\mathbb{P}}({\mathcal{O}}_{Q^{n-1}}\oplus{\mathcal{O}}_{Q^{n-1}}(1))$ and by $\psi:X\to Q^{n-1}$ the natural projection. We have the following exact sequence
$0\to T_{\psi}\to T_{X}\to\psi^{\ast}T_{Q^{n-1}}\to 0$
By [29, II  Exercises 5.16], the above sequence implies the following exact sequence
(1) $0\to T_{\psi}\otimes\wedge^{2}\psi^{\ast}T_{Q^{n-1}}\to\wedge^{3}T_{X}\to\wedge^{3}\psi^{\ast}T_{Q^{n-1}}\to 0$
On the other hand, we have
$T_{\psi}\otimes\wedge^{2}\psi^{\ast}T_{Q^{n-1}}\cong{\mathcal{O}}_{X}(2)\otimes\psi^{\ast}\left\{\left(\wedge^{2}T_{Q^{n-1}}\right)\otimes{\mathcal{O}}_{Q^{n-1}}(-1)\right\}.$
From the standard exact sequence $0\to T_{Q^{n-1}}\to T_{{\mathbb{P}}^{n}}|_{Q^{n-1}}\to{\mathcal{O}}_{Q^{n-1}}(2)\to 0$ and [29, II  Exercises 5.16], we obtain a surjection
$\wedge^{3}T_{{\mathbb{P}}^{n}}|_{Q^{n-1}}\to\left(\wedge^{2}T_{Q^{n-1}}\right)\otimes{\mathcal{O}}_{Q^{n-1}}(2).$
Tensoring ${\mathcal{O}}_{Q^{n-1}}(-3)$ with the above surjection, we see that $\left(\wedge^{2}T_{Q^{n-1}}\right)\otimes{\mathcal{O}}_{Q^{n-1}}(-1)$ is nef. Then (1) implies that $\wedge^{3}T_{X}$ is nef.
Example 4.3.
Let $X={\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{n-1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{n-1}}(-2))$ for $n\geq 4$. Then by the same argument as in Example 4.2 we see that $\wedge^{3}T_{X}$ is nef.
Example 4.4.
Let $X$ be the blow up of ${\mathbb{P}}^{n}$ along a line $\ell$, that is, $X={\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{n-2}}^{\oplus 2}\oplus{\mathcal{O}}_{{\mathbb{P}}^{n-2}}(1))$ for $n\geq 4$. We denote by ${\mathcal{O}}_{X}(1)$ the tautological line bundle of $X={\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{n-2}}^{\oplus 2}\oplus{\mathcal{O}}_{{\mathbb{P}}^{n-2}}(1))$ and by $\psi:X\to{\mathbb{P}}^{n-2}$ the natural projection. We have the following exact sequence
$0\to T_{\psi}\to T_{X}\to\psi^{\ast}T_{{\mathbb{P}}^{n-2}}\to 0$
By [29, II  Exercises 5.16], there exists a vector bundle $F$ on $X$ fitting in the following exact sequences
(2) $0\to F\to\wedge^{3}T_{X}\to\wedge^{3}\psi^{\ast}T_{{\mathbb{P}}^{n-2}}\to 0,$
(3) $0\to{\mathcal{O}}_{X}(3)\otimes\psi^{\ast}\left(T_{{\mathbb{P}}^{n-2}}(-1)\right)\to F\to T_{\psi}\otimes\wedge^{2}\psi^{\ast}T_{{\mathbb{P}}^{n-2}}\to 0$
On the other hand, tensoring the relative Euler sequence with $\wedge^{2}\psi^{\ast}T_{{\mathbb{P}}^{n-2}}$, we obtain a surjection
${\mathcal{O}}_{X}(1)\otimes\psi^{\ast}\left(\left(\wedge^{2}T_{{\mathbb{P}}^{n-2}}\right)^{\oplus 2}\oplus\left(\wedge^{2}T_{{\mathbb{P}}^{n-2}}\right)(-1)\right)\to T_{\psi}\otimes\wedge^{2}\psi^{\ast}T_{{\mathbb{P}}^{n-2}}.$
This implies that $T_{\psi}\otimes\wedge^{2}\psi^{\ast}T_{{\mathbb{P}}^{n-2}}$ is nef. Since ${\mathcal{O}}_{X}(3)\otimes\psi^{\ast}\left(T_{{\mathbb{P}}^{n-2}}(-1)\right)$ is also nef, (3) yields that $\wedge^{3}T_{X}$ is nef.
Example 4.5.
Let $X$ be the blow up of $Q^{n}\subset{\mathbb{P}}^{n+1}$ along a line $\ell$ for $n\geq 4$. Consider the projection of $Q^{n}$ from $\ell$; we denote it by $\pi:Q^{n}\dashrightarrow{\mathbb{P}}^{n-1}$. We can eliminate the point of indeterminacy of $\pi$ by taking the blow up of $Q^{n}$ along $\ell$; then we have the following commutative diagram:
${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}$$\scriptstyle{\psi}$${Q^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${{\mathbb{P}}^{n-1}}$
For any $p\in{\mathbb{P}}^{n-1}$, there exists a plane $\Lambda\subset{\mathbb{P}}^{n+1}$ containing $\ell$ such that the fiber $\pi^{-1}(p)=\Lambda\cap Q^{n}\setminus\ell$. It turns out that $\psi$ is a ${\mathbb{P}}^{1}$-fibration which is not equidimensional. By Theorem 1.2, $\wedge^{3}T_{X}$ is not nef if $n>4$. Thus we assume that $n=4$. In this case, we can prove that $\wedge^{3}T_{X}$ is nef, but we omit the proof because it is routine work (see [73, Example 1.6]). On the other hand, we have
$-K_{X}=\varphi^{\ast}(-K_{Q^{4}})-2E=2\left(2\varphi^{\ast}H_{Q^{4}}-E\right),$
where $E$ is the exceptional divisor of $\varphi$ and $H_{Q^{4}}$ is a hyperplane section of $Q^{4}$. This implies that the Fano index of $X$ is two. This yields that $\tau_{X}\geq 2=n-2$.
Example 4.6.
Let $X$ be the blow up of $Q^{n}\subset{\mathbb{P}}^{n+1}$ along a conic which is not on a plane contained in $Q^{n}$ for $n\geq 4$. By the same argument as in Example 4.5, we see that $\wedge^{3}T_{X}$ is not nef if $n>4$. If $n=4$, we see that $\wedge^{3}T_{X}$ is nef (see [73, Example 1.9]) and the Fano index of $X$ is two. This yields that $\tau_{X}\geq 2=n-2$.
Example 4.7.
For $n\geq 4$, let $Y$ be the blow up of ${\mathbb{P}}^{n-1}$ at a point and $X$ be the product of ${\mathbb{P}}^{1}$ and $Y$. We denote by $p_{1}:X\to{\mathbb{P}}^{1}$ and $p_{2}:X\to{\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{n-2}}(1)\oplus{\mathcal{O}}_{{\mathbb{P}}^{n-2}})$ the first and second projections, respectively. Then we see that $\wedge^{3}T_{X}\cong\left(p_{1}^{\ast}T_{{\mathbb{P}}^{1}}\otimes\wedge^{2}{p_{2}}^{\ast}T_{Y}\right)\oplus\wedge^{3}{p_{2}}^{\ast}T_{Y}$ is nef.
In the rest of this section, we always assume the following:
Assumption 4.8.
Let $X$ be a smooth Fano variety of dimension $n\geq 4$ and $\tau_{X}\geq n-2$. Moreover assume that $X$ admits a birational contraction $\varphi:X\to Y$ of an extremal ray $R$. We denote by $E$ the exceptional locus of $\varphi$ and by $F$ a fiber of $\varphi|_{E}:E\to W:=\varphi(E)$.
Lemma 4.9.
$(\dim E,\dim W)=(n-1,0)$ or $(n-1,1)$. In the latter case, we have $\ell(R)=n-2$.
Proof.
Since any rational curve $C$ contracted by $\varphi$ is not free, we have $-K_{X}\cdot C\geq n-2$. Then Proposition 2.1 yields that
$32(n-1)\geq\dim E+\dim F\geq n+\ell(R)-1\geq n+(n-2)-1=2n-$ 3 .
As a consequence, we obtain our assertion.
Lemma 4.10.
There exists an extremal ray $R^{\prime}$ such that $E\cdot R^{\prime}>0$. Moreover every fiber of the contraction of the ray $R^{\prime}$ has dimension at most two.
Proof.
Take a rational curve $C_{0}\subset X$ such that $E\cdot C_{0}>0$. According to the Cone theorem, there exist finitely many extremal rational curves $C_{i}$’s and positive real numbers $a_{i}$’s such that $[C_{0}]$ is numerically equivalent to $\sum a_{i}[C_{i}]$; since we have
$\sum a_{i}(E\cdot C_{i})=E\cdot C_{0}>0,$
there exists an extremal rational curve $C_{i_{0}}$ such that $E\cdot C_{i_{0}}>0$ as desired.
The latter part of our assertion follows from standard dimension count.
In the following, we denote by $\psi:X\to Y^{\prime}$ the contraction associated to $R^{\prime}$.
Proposition 4.11.
Assume that $\psi$ is of fiber type. If $(\dim E,\dim W)=(n-1,0)$, then $X$ is isomorphic to one of the following:
  1. (i)
    ${\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{n-1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{n-1}}(-1))$.
  2. (ii)
    ${\mathbb{P}}({\mathcal{O}}_{Q^{n-1}}\oplus{\mathcal{O}}_{Q^{n-1}}(-1))$.
  3. (iii)
    ${\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{n-1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{n-1}}(-2))$.
Proof.
We claim that $\psi(E)$ coincides with $Y^{\prime}$. If not, there exists a curve $D$ contained in the exceptional divisor $E$ which is contracted by $\psi$; this is a contradiction, because the curve $[D]$ lives in the rays $R$ and $R^{\prime}$. Then it turns out that $\psi^{-1}(y^{\prime})\cap E$ is a non-empty finite set for any point $y^{\prime}\in Y^{\prime}$. This yields an inequality
$10=\dim\psi^{-1}(y^{\prime})\cap E\geq\dim\psi^{-1}(y^{\prime})+\dim E-n=\dim\psi^{-1}(y^{\prime})-$ 1 .
This tells us that $\dim\psi^{-1}(y^{\prime})\leq 1$ for any point $y^{\prime}\in Y^{\prime}$; since $\psi$ is of fiber type, it is a conic bundle. By our assumption $\tau_{X}\geq n-2$, the anticanonical degree of any rational curve is at least two; this implies that $\psi$ is a ${\mathbb{P}}^{1}$-bundle. Applying [24, Proposition 2.2] and its proof, $Y^{\prime}$ is an $(n-1)$-dimensional smooth Fano variety of $\rho_{Y^{\prime}}=1$ whose Fano index $i_{Y^{\prime}}$ satisfies with $i_{Y^{\prime}}>s>0$; moreover, denoting by ${\mathcal{O}}_{Y^{\prime}}(1)$ the ample generator of ${\rm Pic}(Y^{\prime})$, $X$ is isomorphic to ${\mathbb{P}}({\mathcal{O}}_{Y^{\prime}}\oplus{\mathcal{O}}_{Y^{\prime}}(s))$ and $E$ gives a section of $\psi:X\to Y^{\prime}$.
Let $f$ be a minimal extremal rational curve of the ray $R$; by our assumption $\tau_{X}\geq n-2$ and [18], the curve $f$ satisfies with $-K_{X}\cdot f\geq n-2$ and $n\geq-K_{E}\cdot f$. We have $K_{X}=\varphi^{\ast}K_{Y}+\alpha E$ for some $\alpha\in{\mathbb{Q}}$; then $-\alpha E=-K_{X}+\varphi^{\ast}K_{Y}$ is $\varphi$-nef. Since $\varphi_{\ast}(-\alpha E)=0$ is effective, the negativity lemma ([40, Lemma 3.39]) yields that $-\alpha E$ is effective; then it turns out that $E\cdot f$ is negative. Thus we have
$n\geq-K_{E}\cdot f=-K_{X}\cdot f-E\cdot f\geq(n-2)+1=n-1.$
This implies that $(-K_{E}\cdot f,-K_{X}\cdot f,-E\cdot f)=(n,n-1,1),(n-1,n-2,1)$ or $(n,n-2,2)$.
Let us assume that $(-K_{E}\cdot f,-K_{X}\cdot f,-E\cdot f)=(n,n-1,1)$. Since $Y^{\prime}$ is an $(n-1)$-dimensional smooth Fano variety of $\rho_{Y^{\prime}}=1$, we see that $i_{Y^{\prime}}=n$ and $s=1$; by Kobayashi-Ochiai Theorem [37], we obtain $E\cong Y^{\prime}\cong{\mathbb{P}}^{n-1}$ and $X\cong{\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{n-1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{n-1}}(-1))$. If $(-K_{E}\cdot f,-K_{X}\cdot f,-E\cdot f)=(n-1,n-2,1)$, the same argument implies that $X\cong{\mathbb{P}}({\mathcal{O}}_{Q^{n-1}}\oplus{\mathcal{O}}_{Q^{n-1}}(-1))$. Thus assume that $(-K_{E}\cdot f,-K_{X}\cdot f,-E\cdot f)=(n,n-2,2)$. Applying [18], we see that $E\cong Y^{\prime}\cong{\mathbb{P}}^{n-1}$ and $X\cong{\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{n-1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{n-1}}(-2))$.
Proposition 4.12.
Assume that $\psi$ is of fiber type. If $(\dim E,\dim W)=(n-1,1)$, then $X$ is isomorphic to one of the following:
  1. (i)
    the blow up of ${\mathbb{P}}^{n}$ along a line;
  2. (ii)
    the blow up of $Q^{4}$ along a line;
  3. (iii)
    the blow up of $Q^{4}$ along a conic which is not on a plane contained in $Q^{4}$;
  4. (iv)
    ${\mathbb{P}}^{1}\times{\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{n-2}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{n-2}}(-1))$.
Proof.
By Lemma 4.9, we have $\ell(R)=n-2$; thus [3, Theorem 5.1] implies that $\varphi:X\to Y$ is the blow-up of $Y$ along a smooth curve $W\subset Y$; by our assumption $\tau_{X}\geq n-2$ and [38, II Lemma 3.13], we have $\iota_{X}\geq 2$. Since any fiber $F\cong{\mathbb{P}}^{n-2}$ of $E={\mathbb{P}}(N_{W/Y}^{\vee})\to W$ is not contracted by $\psi$, $\psi|_{F}:F\cong{\mathbb{P}}^{n-2}\to\psi(F)$ is a finite morphism. Thus it turns out that $n-2=\dim\psi(F)\leq\dim Y^{\prime}<n$. Then our assertion follows from [64, Lemma 2, Proposition 3 and Proposition 4], Example 4.5 and Example 4.6. Remark that the pseudoindices of the Fano varieties $3,4$, and $5$ in [64, Proposition 4] are all one. Thus our assertion holds.
Proposition 4.13.
The contraction $\psi:X\to Y^{\prime}$ of the extremal ray $R^{\prime}$ is of fiber type.
Proof.
First assume that there exists an extremal ray $R^{\prime\prime}$ such that $E\cdot R^{\prime\prime}>0$ and the contraction associated to $R^{\prime\prime}$ is of fiber type. Then Proposition 4.11 and 4.12 imply that there is no birational contraction of an extremal ray other than $\varphi$. Hence our assertion holds. As a consequence, we may assume that $\psi$ is of birational type. We denote by $E^{\prime}$ the exceptional divisor of $\psi$ and by $F^{\prime}$ a fiber of $\psi|_{E^{\prime}}:E^{\prime}\to W^{\prime}:=\psi(E^{\prime})$. By the same argument as in Lemma 4.9, we see that $(\dim E^{\prime},\dim\psi(E))=(n-1,0)$ or $(n-1,1)$. By our assumption, $E\cdot R^{\prime}>0$; this implies that $E\cap E^{\prime}\neq\emptyset$. Then we may find a fiber $F_{0}$ of $\varphi|_{E}:E\to W$ and a fiber $F_{0}^{\prime}$ of $\psi|_{E^{\prime}}:E^{\prime}\to W^{\prime}$ such that $F_{0}\cap F_{0}^{\prime}\neq\emptyset$. Then we have
$\dim F_{0}\cap F_{0}^{\prime}\geq\dim F_{0}+\dim F_{0}^{\prime}-\dim X.$
Remark that $\dim F_{0}$ and $\dim F_{0}^{\prime}$ are $n-1$ or $n-2$. If $\dim F_{0}$ or $\dim F_{0}^{\prime}$ is $n-1$, then the above inequality tells us that $\dim F_{0}\cap F_{0}^{\prime}\geq 1$, because we assume that $n\geq 4$. This contradicts to $\varphi\neq\psi$. This yields that $\dim F_{0}=\dim F_{0}^{\prime}=n-2$. By the same argument, we see that the only possible case is $(n,\dim F_{0},\dim F_{0}^{\prime})=(4,2,2)$. We assume this from now on. By Lemma 4.9, we have $\ell(R^{\prime})=2$; thus [3, Theorem 5.1] implies that $\psi:X\to Y^{\prime}$ is the blow-up of $Y^{\prime}$ along a smooth curve $W^{\prime}\subset Y^{\prime}$. Then by [63, Theorem 1], $X$ is isomorphic to the blow-up of ${\mathbb{P}}^{4}$ along a smooth complete intersection of a hyperplane and two quadrics. Then another contraction of $X$ is the blow-up of a singular variety (see [63, 2 An example]). This contradicts to the fact that $Y$ and $Y^{\prime}$ are smooth.
As a consequence, we obtain Theorem 1.3.
Remark 4.14.
In the paper [73], without any argument as Proposition 4.13, Yasutake claim that $\psi$ is of fiber type. This part of his proof is not clear.
5. Fano fourfolds of $\tau_{X}\geq 2$ and $\rho_{X}\geq 2$
In this section, we give a complete classification of smooth Fano fourfolds of $\tau_{X}\geq 2$ and $\rho_{X}\geq 2$. Let us first remark the following:
Lemma 5.1.
For a smooth Fano fourfold of $\rho_{X}\geq 2$, $\tau_{X}\geq 2$ if and only if $\iota_{X}\geq 2$.
Proof.
The “if” part is trivial, and the “only if” part is immediate from [38, II Lemma 3.13].
According to Theorem 1.3, we only need to deal with the case when any elementary contraction is of fiber type. Moreover if $X$ admits a contraction onto ${\mathbb{P}}^{1}$, Proposition 3.1 implies that $X$ is isomorphic to the product of ${\mathbb{P}}^{1}$ and a Fano threefold $Z$ of $\iota_{Z}\geq 2$. Thus, throughout this section, we always assume the following:
Assumption 5.2.
Let $X$ be a smooth Fano fourfold of $\tau_{X}\geq 2$ and $\rho_{X}\geq 2$. Assume that any elementary contraction of $X$ is of fiber type and $X$ does not admit a contraction onto ${\mathbb{P}}^{1}$.
Proposition 5.3.
Under the Assumption 5.2, one of the following holds:
  1. (i)
    $X$ admits a ${\mathbb{P}}^{d}$-bundle structure ($d=1,2$);
  2. (ii)
    $X$ admits two ${\mathbb{P}}^{1}$-fibrations whose special fibers are two-dimensional;
  3. (iii)
    $X$ admits an ${\mathbb{P}}^{1}$-fibration whose special fibers are two-dimensional and an equidimensional $Q^{2}$-fibration;
  4. (iv)
    $X$ admits two equidimensional $Q^{2}$-fibrations.
Proof.
Let us first assume that there exists a smooth elementary contraction $\varphi:X\to Y$. Let $F$ be a fiber of $\varphi$; by [12, Example 3.8], we see that the the Picard number $\rho_{F}$ is one; thus $F$ is a Fano variety of $\rho_{F}=1$. If $\dim F=1$ or $2$, then $\varphi:X\to Y$ is a ${\mathbb{P}}^{d}$-bundle ($d=1,2$). So we assume that $\dim F=3$; in this case, $Y$ is isomorphic to ${\mathbb{P}}^{1}$; this is a contradiction. So we may assume that any elementary contraction is not smooth. Then our assertion follows from Theorem 1.2.
Example 5.4.
Let ${\mathcal{N}}$ be the null correlation bundle on ${\mathbb{P}}^{3}$, which is defined as the kernel of a surjective map $T_{{\mathbb{P}}^{3}}(-1)\to{\mathcal{O}}_{{\mathbb{P}}^{3}}(1)$ (see for instance [52, Section 4.2]). Let ${\mathcal{S}}$ be the spinor bundle on $Q^{3}$, which is by definition the restriction to $Q^{3}$ of the universal quotient bundle of the Grassmannian $G(1,{\mathbb{P}}^{3})$ (see for instance [53]). Then ${\mathbb{P}}({\mathcal{N}})$ is isomorphic to ${\mathbb{P}}({\mathcal{S}})$ and we have the following diagram:
[…]
The variety ${\mathbb{P}}({\mathcal{N}})\cong{\mathbb{P}}({\mathcal{S}})$ is a Fano fourfold of index two.
Example 5.5.
We continue to follow the notation in Example 5.4. If we take a suitable double cover $\pi:Q^{3}\to{\mathbb{P}}^{3}$ whose branch divisor $B$ is a quadric, then $X={\mathbb{P}}(\pi^{\ast}{\mathcal{N}})$ is a Fano fourfold of index one (see [60, Corollary]). Remark that $X\cong Q^{3}\times_{{\mathbb{P}}^{3}}{\mathbb{P}}({\mathcal{N}})$ satisfying the following commutative diagram:
${X\cong Q^{3}\times_{{\mathbb{P}}^{3}}{\mathbb{P}}({\mathcal{N}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{q_{1}}$$\scriptstyle{q_{2}}$$\scriptstyle{f}$${Q^{3}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${{\mathbb{P}}({\mathcal{N}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p_{1}}$${{\mathbb{P}}^{3}}$
where $q_{1}:X\to Q^{3}$ and $q_{2}:X\to{\mathbb{P}}({\mathcal{N}})$ are natural projections and $f$ is the composition $\pi\circ q_{1}$. We prove that $\iota_{X}=1$. To prove this, let $\ell$ be a line on $Q^{3}$ such that $\pi(\ell)\subset{\mathbb{P}}^{3}$ is also a line. By replacing $\ell$ if necessary, we may assume that there exists a fiber $\ell_{2}$ of $p_{2}$ which is isomorphic to $\pi(\ell)$ via $p_{1}$. For a rational curve $\ell_{X}:=\ell\times_{{\mathbb{P}}^{3}}\ell_{2}\hookrightarrow X$,
$-K_{X}\cdot\ell_{X}$ $=$ $f^{\ast}K_{{\mathbb{P}}^{3}}\cdot\ell_{X}+{q_{1}}^{\ast}(-K_{Q^{3}})\cdot\ell_{X}+{q_{2}}^{\ast}(-K_{{\mathbb{P}}({\mathcal{N}})})\cdot\ell_{X}$
$=$ $K_{{\mathbb{P}}^{3}}\cdot\pi(\ell)+(-K_{Q^{3}})\cdot\ell+(-K_{{\mathbb{P}}({\mathcal{N}})})\cdot\ell_{2}$
$=$ $-4+3+2=1$
Thus we see that $\iota_{X}=1$.
Proposition 5.6.
Under the assumption of 5.2, assume that $X$ admits a ${\mathbb{P}}^{1}$-bundle structure $\varphi:X\to Y$. Then $X$ is isomorphic to ${\mathbb{P}}({\mathcal{N}})$, where ${\mathcal{N}}$ is the null correlation bundle over ${\mathbb{P}}^{3}$.
Proof.
Let us first claim that $\rho_{X}=2$. To prove this, assume the contrary, that is, $\rho_{X}\geq 3$. Then it follows from [39, Corollary 2.9] that $Y$ is a smooth Fano threefold of $\rho_{Y}\geq 2$. Applying [23, Lemma 4.2, Lemma 4.4 and Theorem 6.3] (remark that ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ is missing in the list of [23, Theorem 6.3]), we see that $Y$ is one of the following:
  1. (1)
    A double cover of $W={\mathbb{P}}(T_{{\mathbb{P}}^{2}})$ whose branch locus is a divisor in $|-K_{W}|$.
  2. (2)
    A divisor on ${\mathbb{P}}^{2}\times{\mathbb{P}}^{2}$ of bidegree $(1,1)$, which is isomorphic to ${\mathbb{P}}(T_{{\mathbb{P}}^{2}})$.
  3. (3)
    A divisor on ${\mathbb{P}}^{2}\times{\mathbb{P}}^{2}$ of bidegree $(1,2)$.
  4. (4)
    A divisor on ${\mathbb{P}}^{2}\times{\mathbb{P}}^{2}$ of bidegree $(2,2)$.
On the other hand, it follows from [6, Lemma 2.5] that $\iota_{Y}\geq\iota_{X}\geq 2$. Since the pseudoindex of the Fano variety appearing in $(1)$, $(3)$, and $(4)$ is one (for variety in $(1)$, see [46, p. 53, $n^{o}$ 6 (6.B)]), we are in the case $(2)$, that is, $Y$ is isomorphic to ${\mathbb{P}}(T_{{\mathbb{P}}^{2}})$. Let $Y$ be ${\mathbb{P}}(T_{{\mathbb{P}}^{2}})$. Since $Y$ is rational, its Brauer group vanishes. This yields that there exists a rank two vector bundle ${\mathcal{E}}$ on $Y$ such that $X={\mathbb{P}}({\mathcal{E}})$ and $\varphi$ is the natural projection $X={\mathbb{P}}({\mathcal{E}})\to Y$. Remark that $\iota_{X}=\iota_{Y}=2$. Since ${\mathbb{P}}(T_{{\mathbb{P}}^{2}})$ is isomorphic to a divisor on ${\mathbb{P}}^{2}\times{\mathbb{P}}^{2}$ of bidegree $(1,1)$, there exist two projections $p_{i}:{\mathbb{P}}(T_{{\mathbb{P}}^{2}})\to{\mathbb{P}}^{2}$ for $i=1,2$ as follows:
${{\mathbb{P}}(T_{{\mathbb{P}}^{2}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p_{1}}$$\scriptstyle{p_{2}}$${{\mathbb{P}}^{2}}$${{\mathbb{P}}^{2}}$
In this diagram, the projective plane ${\mathbb{P}}^{2}$ on the left can be seen as a unique family of lines on the projective plane ${\mathbb{P}}^{2}$ on the right; then this diagram is nothing but the universal family of lines. For any fiber $\ell_{i}$ of $p_{i}$, [6, Lemma 2.5] tells us that ${\mathcal{E}}|_{\ell_{i}}\cong{\mathcal{O}}_{{\mathbb{P}}^{1}}(a_{i})^{\oplus 2}$ for some $a_{i}\in{\mathbb{Z}}$ $(i=1,2)$. Tensoring a line bundle, we may assume that ${\mathcal{E}}|_{\ell_{i}}\cong{\mathcal{O}}_{{\mathbb{P}}^{1}}^{\oplus 2}$ for any fiber $\ell_{i}$ of $p_{i}$. Then it follows from Grauert’s theorem and Nakayama’s lemma that ${p_{2}}_{\ast}{\mathcal{E}}$ is a rank $2$ vector bundle on ${\mathbb{P}}^{2}$ and ${p_{2}}^{\ast}({p_{2}}_{\ast}{\mathcal{E}})\cong{\mathcal{E}}$. Moreover we have
$({p_{2}}_{\ast}{\mathcal{E}})|_{p_{2}(\ell_{1})}\cong{p_{2}}^{\ast}({p_{2}}_{\ast}{\mathcal{E}})|_{\ell_{1}}\cong{\mathcal{E}}|_{\ell_{1}}\cong{\mathcal{O}}_{{\mathbb{P}}^{1}}^{\oplus 2}.$
This means that ${p_{2}}_{\ast}{\mathcal{E}}$ is a uniform vector bundle on ${\mathbb{P}}^{2}$; then [65] implies that ${p_{2}}_{\ast}{\mathcal{E}}\cong{\mathcal{O}}_{{\mathbb{P}}^{2}}^{\oplus 2}$. As a consequence, $X$ is isomorphic to ${\mathbb{P}}^{1}\times{\mathbb{P}}(T_{{\mathbb{P}}^{2}})$, which contradicts to our assumption. Thus we see that $\rho_{X}=2$.
Since $Y$ is a Fano threefold of $\rho_{Y}=1$ and $\iota_{Y}\geq 2$, [50, Lemma 2.16] implies that $Y$ is ${\mathbb{P}}^{3}$ or $Q^{3}$. Since $Y$ is rational, there exists a rank two vector bundle ${\mathcal{E}}$ on $Y$ such that $X={\mathbb{P}}({\mathcal{E}})$. From the classification of rank two Fano bundles over ${\mathbb{P}}^{3}$ [61, Theorem 2.1] and that of $Q^{3}$ [59, Theorem], $X$ is isomorphic to one of the following:
  1. (i)
    ${\mathbb{P}}({\mathcal{N}})$, where ${\mathcal{N}}$ is the null correlation bundle over ${\mathbb{P}}^{3}$;
  2. (ii)
    ${\mathbb{P}}({\mathcal{S}})$, where ${\mathcal{S}}$ is the spinor bundle over $Q^{3}$;
  3. (iii)
    ${\mathbb{P}}({\pi}^{\ast}{\mathcal{N}})$, where ${\mathcal{N}}$ is the null correlation bundle over ${\mathbb{P}}^{3}$ and $\pi:Q^{3}\to{\mathbb{P}}^{3}$ is a double cover.
By Example 5.4 and Example 5.5, $X$ is isomorphic to ${\mathbb{P}}({\mathcal{N}})\cong{\mathbb{P}}({\mathcal{S}})$.
Proposition 5.7.
Under the assumption of 5.2, assume that $X$ admits a ${\mathbb{P}}^{2}$-bundle structure $\varphi:X\to Y$. Then $X$ is isomorphic to one of the following:
  1. (i)
    ${\mathbb{P}}^{2}\times{\mathbb{P}}^{2}$;
  2. (ii)
    a divisor on ${\mathbb{P}}^{2}\times{\mathbb{P}}^{3}$ of bidegree $(1,1)$;
Proof.
By [39, Corollary 2.9] and [23, Lemma 4.2, Lemma 4.4 and Proposition 6.1], we see that $Y$ is isomorphic to ${\mathbb{P}}^{2}$ or ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$. In the latter case, $X$ admits a contraction onto ${\mathbb{P}}^{1}$, which contradicts to our assumption. Thus $Y$ is isomorphic to ${\mathbb{P}}^{2}$; then our assertion follows from [62].
Proposition 5.8.
Under the assumption of 5.2, assume that $X$ satisfies one of the following:
  1. (i)
    $X$ admits two ${\mathbb{P}}^{1}$-fibrations whose special fibers are two-dimensional;
  2. (ii)
    $X$ admits a ${\mathbb{P}}^{1}$-fibration whose special fibers are two-dimensional and an equidimensional $Q^{2}$-fibration.
Then $\rho_{X}=i_{X}=2$.
Proof.
In both cases, there exists a ${\mathbb{P}}^{1}$-fibration whose special fibers are two-dimensional, which is denoted by $\varphi:X\to Y$. Remark that the structure of $\varphi$ is described in [34]. By [34, Lemma 5.2], there exist only finitely many $2$-dimensional fibers of $\varphi$, which are denoted by $F_{1}=\varphi^{-1}(y_{1}),\ldots,F_{m}=\varphi^{-1}(y_{m})$. By [34, Theorem 0.6] (see also [4, Lemma 5.9.4]), $Y$ is smooth and we have $F_{i}\cong{\mathbb{P}}^{2}$ and $N_{F_{i}/X}\cong\Omega_{{\mathbb{P}}^{2}}(1)$. Let $\pi:\tilde{X}\to X$ be the blow-up of $X$ along $F_{1}\sqcup\ldots\sqcup F_{m}$ and $\tilde{Y}$ the blow-up of $Y$ along $\{y_{1},\ldots,y_{m}\}$. Then the exceptional divisor $E$ of $\tilde{X}\to X$ is
${\mathbb{P}}(N_{F_{1}/X}^{\vee})\sqcup\ldots\sqcup{\mathbb{P}}(N_{F_{m}/X}^{\vee})\cong{\mathbb{P}}(T_{{\mathbb{P}}^{2}}(-1))\sqcup\ldots\sqcup{\mathbb{P}}(T_{{\mathbb{P}}^{2}}(-1))$
By the universal property of blow-up (see for instance [29, Proposition 7.14, Corollary 7.15]), there exists a morphism $\tilde{\varphi}:\tilde{X}\to\tilde{Y}$
[…]
making a commutative diagram as shown. Remark that $\tilde{\varphi}:\tilde{X}\to\tilde{Y}$ is a ${\mathbb{P}}^{1}$-bundle (cf. [2, Remark 4.13]). Each irreducible component ${\mathbb{P}}(N_{F_{i}/X})\cong{\mathbb{P}}(T_{{\mathbb{P}}^{2}}(-1))$ of the exceptional divisor $E$ is isomorphic to a hyperplane section of ${\mathbb{P}}^{2}\times{\mathbb{P}}^{2}$; the restrictions $\tilde{\varphi}|_{{\mathbb{P}}(N_{F_{i}/X}^{\vee})}$ and $\pi|_{{\mathbb{P}}(N_{F_{i}/X}^{\vee})}$ give two natural projections of ${\mathbb{P}}(T_{{\mathbb{P}}^{2}}(-1))\subset{\mathbb{P}}^{2}\times{\mathbb{P}}^{2}$ onto ${\mathbb{P}}^{2}$.
By our assumption, we have another elementary contraction $\psi:X\to Z$ which is
  • a ${\mathbb{P}}^{1}$-fibration whose special fibers are two-dimensional; or
  • an equidimensional $Q^{2}$-fibration.
In each case, $Z$ is a smooth variety of dimension at most three (see [34, Theorem 0.6] and [1, Proposition 1.4.1]). We proceed as the proof of [12, Theorem 4.1]. Now we have the following diagram:
[…]
This yields a morphism $\tilde{Y}\to{\rm Chow}(Z)$; we denote by $Y^{\prime}$ the image of this morphism. Let us consider a geometric quotient of $Z$ for $Y^{\prime}$; by applying [7, Corollary 1], there exists an elementary contraction $f:Z\to W$ which contracts a rational curve $\psi\circ\pi(\tilde{\varphi}^{-1}(\tilde{y}))$ for any closed point $\tilde{y}\in\tilde{Y}$. This yields that the surface $\psi(F_{i})$ is contracted by $f$. Since $f$ is constructed as a geometric quotient for $Y^{\prime}$, it is of fiber type; if the relative dimension of $f$ is one, then $f$ is a conic bundle, which contradicts to the fact that $f$ contracts a surface. This yields that $f$ has relative dimension at least two. However if the relative dimension of $f$ is two, then $W$ is isomorphic to ${\mathbb{P}}^{1}$; it turns out that $X$ admits a contraction $f\circ\psi:X\to{\mathbb{P}}^{1}$; this contradicts to our assumption. As a consequence, we see that $W$ is a point and $\rho_{Z}=1$. This implies that $\rho_{X}=2$. Since $Z$ is a uniruled variety of $\rho_{Z}=1$, it is a Fano variety. Since $\dim Z=2$ or $3$, $Z$ is ${\mathbb{P}}^{2}$ or a Fano threefold of $\rho_{Z}=1$. Then there exists a rational curve $\ell\subset Z$ such that $-K_{Z}\cdot\ell$ is equal to the Fano index of $Z$ (see for instance [66]).
Claim 5.9.
There exists a rational curve ~$\tilde{\ell}\subset X$ such that $\psi|_{\tilde{\ell}}:\tilde{\ell}\to\ell$ is birational.
If $\psi$ is a ${\mathbb{P}}^{1}$-fibration which is not equidimensional, then it is a smooth morphism away from finitely many fibers (see [34, Lemma 5.2]). This means that $\psi|_{\psi^{-1}({\ell})}:\psi^{-1}({\ell})\to\ell$ is smooth over a nonempty open subset of $\ell$. Thus the existence of a desired rational curve ~$\tilde{\ell}\subset X$ follows from the same argument as in [67, Section 4.3]. So assume that $\psi:X\to Z$ is an equidimensional $Q^{2}$-fibration. In this case, $Z$ is isomorphic to ${\mathbb{P}}^{2}$; then, by replacing $\ell$ if necessary, we may assume that $\psi|_{\psi^{-1}({\ell})}:\psi^{-1}({\ell})\to\ell$ is smooth over a nonempty open subset of $\ell$. Thus we are done.
By this claim, we may take a family of minimal sections ${\mathcal{M}}\subset{\rm RatCurves}^{n}(X)$ as in [67, Section 4.3], which is unsplit. Moreover we claim the following:
Claim 5.10.
Any ${\mathcal{M}}$-curve is contracted by $\varphi$. In particular, $\tilde{\ell}$ is contracted by $\varphi$.
To prove this, assume the contrary, that is, there exists an ${\mathcal{M}}$-curve which is not contracted by $\varphi$. Then any ${\mathcal{M}}$-curve is not contracted by $\varphi$. If ${\mathcal{M}}$ is a covering family, it follows from [7, Corollary 1] that $\varphi:X\to Y$ is the geometric quotient for ${\mathcal{M}}$. This contradicts to our assumption. Thus $\mathop{\rm codim}\nolimits_{X}{\rm Locus}({\mathcal{M}})\geq 1$; then by Proposition 2.4, for any point $x\in{\rm Locus}({\mathcal{M}})$, we have
$\dim{\rm Locus}({\mathcal{M}}_{x})\geq\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}+\mathop{\rm codim}\nolimits_{X}{\rm Locus}({\mathcal{M}})-1.$
By our assumption, we have $\mathop{\rm deg}\nolimits_{(-K_{X})}{\mathcal{M}}\geq 2$; combining this with the above inequality, we see that $\dim{\rm Locus}({\mathcal{M}}_{x})\geq 2$ for any point $x\in{\rm Locus}({\mathcal{M}})$ and $\mathop{\rm codim}\nolimits_{X}{\rm Locus}({\mathcal{M}})=1$. Let us put $D:={\rm Locus}({\mathcal{M}})$. Here we prove that there exists a curve $\ell_{\varphi}\subset D$ (resp. $\ell_{\psi}\subset D$) which is contracted by $\varphi$ (resp. $\psi$). If $\dim D>\dim\varphi(D)$ and $\dim D>\dim\psi(D)$, then we may find such curves. So assume $\dim D=\dim\varphi(D)=\dim\psi(D)$. Since $\varphi$ and $\psi$ have two-dimensional fibers and $D$ is a divisor on $X$, by dimension count we may find desired curves.
We claim that the curve $\ell_{\varphi}$ is not contained in ${\rm Locus}({\mathcal{M}}_{x})$ for any $x\in D$. In fact, if $\ell_{\varphi}$ is contained in ${\rm Locus}({\mathcal{M}}_{x_{0}})$ for some $x_{0}\in D$, it follows from [38, II Corollary 4.21] that any ${\mathcal{M}}$-curve is contracted by $\varphi$; this is a contradiction. Thus we see that $\dim{\rm Locus}({\mathcal{M}}_{x})=2$ for any $x\in D$ and we obtain
$D=\bigcup_{x\in\ell_{\varphi}}{\rm Locus}({\mathcal{M}}_{x})$
By [51, Lemma 3.2, Remark 3.3], there exist rational numbers $\lambda\geq 0,\mu$ and $[C]\in{\mathcal{M}}$ such that $\ell_{\psi}=\lambda\ell_{\varphi}+\mu C\in N_{1}(X)$. For any ample divisor $A$ on $Z$, we have
$0=\psi^{\ast}(A)\cdot\ell_{\psi}=\lambda\psi^{\ast}(A)\cdot\ell_{\varphi}+\mu\psi^{\ast}(A)\cdot C.$
This implies that $\mu\leq 0$. Therefore we obtain
$\ell_{\psi}+(-\mu)C=\lambda\ell_{\varphi}\in{\mathbb{R}}_{\geq 0}[\ell_{\varphi}].$
Since ${\mathbb{R}}_{\geq 0}[\ell_{\varphi}]$ is an extremal ray, this yields $\ell_{\psi}\in{\mathbb{R}}_{\geq 0}[\ell_{\varphi}]$; this is a contradiction. As a consequence, Claim 5.10 holds.
On the other hand, since $Y$ is a Fano threefold of $\rho_{Y}=1$, we may find a rational curve $\ell^{\prime}\subset Y$ such that $-K_{Y}\cdot\ell^{\prime}=i_{Y}$. By the same argument as in Claim 5.9 and Claim 5.10, we may find a rational curve ~$\tilde{\ell^{\prime}}\subset X$ which is contracted by $\psi$.
Let $H_{Y}$ and $H_{Z}$ be the ample generators of $\rm Pic(Y)$ and $\rm Pic(Z)$ respectively. Since $\rho_{X}=2$, then $\{\varphi^{\ast}H_{Y},\psi^{\ast}H_{Z}\}$ is a basis of $N^{1}(X)$. Thus there exist rational numbers $a,b$ such that $-K_{X}=a\varphi^{\ast}H_{Y}+b\psi^{\ast}H_{Z}$. Denoting by $C_{\varphi}$ and $C_{\psi}$ minimal extremal rational curves of the rays associated to $\varphi$ and $\psi$ respectively, we have
$2=-K_{X}\cdot C_{\psi}=aH_{Y}\cdot\varphi_{\ast}(C_{\psi}).$
On the other hand, we also have
$2\leq-K_{X}\cdot\tilde{\ell^{\prime}}=aH_{Y}\cdot\varphi_{\ast}(\tilde{\ell^{\prime}})=a.$
These imply that $H_{Y}\cdot\varphi_{\ast}(C_{\psi})=1$ and $a=2$. By the same way, we also obtain $H_{X}\cdot\psi_{\ast}(C_{\varphi})=1$ and $b=2$. As a consequence, $-K_{X}=2(\varphi^{\ast}H_{Y}+\psi^{\ast}H_{Z})$; this yields that $i_{X}=2$.
Proposition 5.11.
Under the assumption of 5.2, assume that $X$ admits two equidimensional $Q^{2}$-fibrations. Then $\rho_{X}=i_{X}=2$.
Proof.
This can be proved by the same argument as in Proposition 5.8.
Summing up, we obtain Theorem 1.4.
6. Fano varieties with nef third exterior power of the tangent bundle
In this short section, we study Fano varieties with nef third exterior power of the tangent bundle. Before delving into the results of this section, a comprehensive understanding of the motivation and historical context of the problem will be provided. In a pioneering study, Demailly, Peternell, and Schneider [22] established that a smooth projective variety with a nef tangent bundle becomes a locally trivial Fano fibration over an Abelian variety up to the differences in the finite étale covers. We call this result a decomposition theorem for varieties with nef tangent bundle. Also, [22] and Solá Conde, Wiśniewski [58, Theorem 4.4] show that all Mori contractions are smooth. We call this result a contraction theorem for varieties with nef tangent bundle. After that, similar problems were considered for varieties with nef (strictly nef or ample) second exterior power of the tangent bundles [9, 19, 72, 73, 55, 45], and finally the problem was solved by the author [67] including all previous results. Although a similar problem can be conceived for varieties $X$ where the $r$-th exterior power of the tangent bundle is nef for $r<\dim X$, a decomposition theorem was obtained by the author [68] via Gachet’s result [17]:
Theorem 6.1 ([68]).
Let $X$ be a smooth projective variety of dimension $n$. Assume that $\wedge^{r}T_{X}$ is nef for some $1\leq r<n$. Then if we take a suitable finite étale cover $\tilde{X}\to X$, there exists a locally trivial fibration $\varphi:\tilde{X}\to A$ such that the fiber $F$ is a Fano variety and $A$ is an Abelian variety. Moreover, if $\dim A\geq r-1$, then $T_{X}$ is nef; otherwise $\wedge^{r-\dim A}T_{\tilde{X}/A}$ is nef.
If $\wedge^{r}T_{X}$ is nef, then so is $\wedge^{r+1}T_{X}$ by [42, Theorem 3.3]. Thus Theorem 6.1 is a generalization of [67, Corollary 1.6]. According to this theorem, the study of smooth projective varieties $X$ with nef $\wedge^{r}T_{X}$ $(1\leq r<n)$ is reduced to the case of Fano varieties.
Based on the above background, in the following, we deal with Fano varieties $X$ with nef $\wedge^{3}T_{X}$. The matter warranting our attention is a contraction theorem for Fano varieties $X$ with nef $\wedge^{3}T_{X}$. Let us first remark that the positivity of the exterior power of the tangent bundle and our new invariant $\tau_{X}$ are involved as follows:
Lemma 6.2 ([9, Lemma 1.3], [72, Lemma 2.9], [67, Lemma 3.2]).
Let $X$ be a smooth Fano variety of dimension $n\geq 4$. If $\wedge^{r}T_{X}$ is nef for some $1\leq r<n$, then we have $\tau_{X}\geq\dim X-r+1$.
By this lemma, if $X$ is a smooth Fano variety of dimension $n\geq 4$ such that $\wedge^{3}T_{X}$ is nef, then $\tau_{X}\geq n-2$; thus Theorem 1.2 and Theorem 1.3 can be seen as a generalization of [22, Theorem 5.2] and [67, Theorem 1.7]. Moreover applying Theorem 1.2, Theorem 1.3 and Theorem 6.1, [67, Theorem 1.9 and 1.10] can be generalized to the case when $\wedge^{3}T_{X}$ is nef.
Acknowledgements. The author would like to thank Taku Suzuki for carefully reading and pointing out typos and for helpful suggestions. The author would like to thank J. Wisńiewski for answering the author’s question on Fano fourfolds and sending his paper [69]. He would also like to thank Bruno Dewer for pointing out typos.
Conflict of Interest. The author has no conflicts of interest directly relevant to the content of this article.
Data availability. Data sharing not applicable to this article as no data sets were generated or analyzed during the current study.